OpenStudy (anonymous):

Ignore air resistance. A diver starts 30 feet above the water (about the height of an Olympic platform diver). What is the diver’s velocity at impact?

OpenStudy (anonymous):

@experimentX can u help?

OpenStudy (experimentx):

\[ v^2 = u^2 + 2 a s\] u = 0 (initial vecolity v = final velocity a = acceleration due to gravity (9.8 m/s^2) <-- change it into feet/sec^2 s = 30 feet calculate it

OpenStudy (anonymous):

here is what i ddi so far can u check? 2(9.8) (9.14) = vf^2 - 0^2

OpenStudy (anonymous):

using second equation of motion 2as = vf^2 -vi^2

OpenStudy (anonymous):

putting it as a = g =9.8 s = 30 feet = i changed it in meters vi =0 then to find vf

OpenStudy (anonymous):

ok

OpenStudy (experimentx):

ok ok